Angle A is 80 less than tree times angle B. If placed side by side they form a straight angle, find each angle.
Answer by checkley77(12844)

x+(3x-80)=180
4x=180+80
4x=260
x=260/4
x=65 ans. for the smaller angle.
Proof:
65+(3*65-80)=180
65+(195-80)=180
65+115=180
180=180